Based on the data shown below, calculate the regression line (each value to three decimal places)

Y= _____ X+ ________
(Caution, make sure you have these entered in the correct order)

X Y
1 18.4
2 21.5
3 23.8
4 29.3
5 31.1
6 33.7
7 37.2
8 37.9
9 40.8
10 43
11 46.9
12 51.3
13 52.5
14 56.3
Answer #1

The observation table is

X Y X*Y X2 Y2
1 18.4 18.4 1 338.56
2 21.5 43 4 462.25
3 23.8 71.4 9 566.44
4 29.3 117.2 16 858.49
5 31.1 155.5 25 967.21
6 33.7 202.2 36 1135.69
7 37.2 260.4 49 1383.84
8 37.9 303.2 64 1436.41
9 40.8 367.2 81 1664.64
10 43 430 100 1849
11 46.9 515.9 121 2199.61
12 51.3 615.6 144 2631.69
13 52.5 682.5 169 2756.25
14 56.3 788.2 196 3169.69
Sum = 105 523.7 4570.7 1015 21419.77

Based on the above table105 i = 7.5 14 7l 7t 523.7 i =-14-= 37.4071 ssxx =Σ x?-I(Σ s) = 1015-1052/14-227.5 SSXY = Σ xx-1 ( Σ x. ) ( Σ x, = 1015-1052/14 = 227.5 Y; | 4570.7-105×523.7/14 642.95

Therefore, based on the above calculations, the regression coefficients (the slope m, and the y-intercept a) are obtained as follows:


642.95 Sxx22752.8262 m - SSxy ΓΙ a y-X . m = 37.4071-7.5 × 2.8262 = 16.211

Therefore, we find that the regression equation is:

2.826 * X + 16.211